# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ Abstract
We propose LoRA-MCL, a training scheme that extends next-token prediction in language models with a method designed to decode diverse, plausible sentence continuations at inference time. Traditional language modeling is an intrinsically ill-posed problem: given a context, multiple futures may be equally plausible. Our approach leverages Multiple Choice Learning (MCL) and the Winner-Takes-All (WTA) loss to efficiently handle ambiguity through Low-Rank Adaptation (LoRA). We provide a theoretical interpretation of applying Multiple Choice Learning to Language Modeling, assuming the data is generated from a mixture of distributions. To illustrate the proposed approach, we use data sampled from mixtures of Markov chains. We then demonstrate with extensive experiments on real-world visual and audio captioning tasks that our method achieves high diversity and relevance in generated outputs.
**footnotetext: Equal Contribution.
@@ 1 Introduction
Predicting what a person will say next or describing the content of an audio or visual scene with text is difficult, if not impossible, to do with perfect accuracy. When the context is not informative enough, external factors may lead to different scenarios or modes of plausible text continuations yang2018breaking ; dieng2020topic ; mei2022diverse . In such ambiguous tasks, the conditional distribution over the space of output sentences given the input context may be multi-modal due to the underlying inherent uncertainty malinin2020uncertainty .
@@
Initially proposed for text processing, transformer-based language models have quickly become a general framework for modeling streams of tokens, which can also represent, for instance, images and audio signals touvron2023llama ; chu2024qwen2 ; liu2023visual . Such models are trained as next-token predictors and allow, by nature, for addressing this uncertainty, by generating plausible output sentences through a wide body of maximum a posteriori (MAP) or sampling-based decoding approaches welleckdecoding . While sampling allows exploration and diversity, it may lead to unreliable responses and requires truncation to avoid unexpected answers, partly due to the overestimation of low-probability tokens zhang2021trading ; hewitt2022truncation . When seeking reliable and expected answers, maximum a posteriori estimation techniques, like Beam Search beamSearch , look for sentences that maximize the model’s likelihood. However, these alternatives have drawbacks as they may lack diversity, be prone to repetition loops keskar2019ctrl , and may sound unnatural holtzmancurious . Some approaches, e.g., Diverse Beam Search (DBS) vijayakumar2018diverse , were therefore proposed to artificially increase the diversity at inference, e.g., through a diversity penalty parameter $\lambda$, to find a tradeoff between generation quality and sample diversity sucontrastive . In contrast with these methods, our approach aims to predict diverse sentences reflecting the ambiguity of the input context.
@@
Multiple Choice Learning (MCL) guzman2012multiple ; lee2016stochastic has emerged as a paradigm for addressing ambiguous tasks. It generally consists of a network with a shared backbone and multiple output heads. During training, it utilizes the winner-takes-all loss for adaptively updating the head that performs the best for each example. This is a competitive training scheme that specializes each model to subsets of the conditional output distribution rupprecht2017learning ; letzelter2024winner . In this paper, we propose incorporating this idea for language model finetuning, leveraging multiple Low-Rank Adapters hu2022lora instead of multiple heads, which may be impractical due to computation requirements and architectural constraints. Our method natively generates diverse and plausible sequences in a single forward pass, aiming to best approximate the conditional output distribution.
Our contributions are as follows:
@@
We propose a new paradigm that adapts MCL for token sequence modeling, with a method, LoRA-MCL, that is particularly suited for efficient finetuning of language models.
@@
We provide a theoretical analysis of our approach, which applies MCL to language modeling. Assuming the sequences are sampled from a mixture of distributions, we explain in particular why LoRA-MCL should enable capturing the modes of the data distribution. We then consider the case of Markov chains, for which we validate our claims with a well-designed toy example.
We undertake an extensive experimental study to validate our approach, on audio and machine vision captioning tasks, demonstrating its wide applicability in challenging real-world scenarios and showing an excellent diversity–quality trade-off. The code will be made publicly available.
@@ 2 Problem setup
$\mathcal{L}(\theta)=-\mathbb{E}_{c,x}[\log p_{\theta}(x\,|\,c)]=\mathbb{E}_{c,x}\Big{[}-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)\Big{]}\;,$ (1)
where $x_{<t}$ denotes the sequence of tokens prior to time $t$.
Optimizing (1) is referred to as teacher-forcing williams1989learning , where $p_{\theta}$ is fed with target (instead of predicted) tokens $x_{<t}$ during training sutskever2014sequence . When using a transformer architecture vaswani2017attention , this is implemented with causal attention modules, which allow for computing the conditional distributions in parallel through all time steps within a single forward pass.
During inference, decoding methods welleckdecoding proceed to generating sequences $\hat{x}$ from the trained model $p_{\theta}$ in an auto-regressive fashion. First, they start with a conditional distribution for the first token from the context: $p_{\theta}(x_{1}\,|\,c)$, which allows selecting $\hat{x}_{1}$. Then, for $t\geq 2$ they predict $p_{\theta}(x_{t}\,|\,\hat{x}_{<t},c)$, and select $\hat{x}_{t}$, until reaching either the sequence length limit or an end-of-sentence (EOS) token. The choice of the decoding method used to generate $K$ candidate sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ depends on the purpose of the task, but the general goal is $(i)$ to get highly likely sentences, i.e., ones that maximize $p_{\theta}(\hat{x})$; and $(ii)$ to get sufficiently diverse sentences, as can be measured by $n$-gram similarity ippolito2019comparison .
Although this is a widely adopted paradigm, we show next how this training and decoding pipeline can be improved: instead of artificially generating diversity at inference time, we aim at learning to predict sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ that cover well the modes of the target distribution $p(x\,|\,c)$.
@@ 3.1 Motivation
In language modeling, topic models nigam2000text ; blei2003latent ; dieng2020topic are data-generating processes in which the ground-truth probability distribution of sequences is modeled as a mixture of latent components or topics. For example, the sequence “I am eating …” may have multiple plausible continuations, but the likelihood of each depends heavily on contextual factors such as the speaker’s location, which influences their culinary habits. Each location (or context) can thus be associated with a distinct word distribution. In topic models, data generation proceeds by first sampling a topic $z\in\mathcal{Z}$ for each sentence (usually referred to as a document in the literature), and then sampling words (or $n$-grams) from the distribution associated with that topic.
With this in mind, MLE in (1) may not be suitable yang2018breaking . While MLE is effective for estimating the overall distribution $p(x)$, it does not capture the individual components when $p(x)$ is expressed as a mixture, i.e., $p(x)=\sum_{k}p(z_{k})p(x\,|\,z_{k})$. In such cases, MLE tends to model the aggregate rather than distinguish topic-specific distributions $p(x\,|\,z_{k})$.
3.2 Applying Multiple Choice Learning to language modeling
Our approach is inspired by the multiple choice learning (MCL) literature guzman2012multiple ; lee2016stochastic . We propose the following training scheme, intending to enable the recovery of the different topics $z_{k}$. Instead of a single model, we consider a set of models $(\theta_{1},\dots,\theta_{K})$. Then the objective (1) is replaced by one consisting of iterating between the following two steps:
@@ 3.2 Applying Multiple Choice Learning to language modeling
    For each training sample $(c,x)$ in the batch $\mathcal{B}$: Compute $p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$, and choose the best model $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$.
@@
    $\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big{[}\max_{k=1,\dots,K}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (2)
@@
In practice when using a finite batch size, (2) is approximated with
$\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})\simeq-\sum_{k=1}^{K}\sum_{\{(x,c)\in\mathcal{B}\;\,|\,\;k^{\star}(x,c)=k\}}\log p(x\,|\,c;\theta_{k})\;,$ (3)
where each hypothesis $k$ is trained only on the subset of examples in which it is the winner, according to the assignment $k^{\star}(x,c)$.
This training procedure, similar to a hard-EM style optimization min2019discrete ; wen2023an , is a competitive training scheme that encourages the different models to explore different areas of the data distribution. However, it is subject to two main issues: First, using $K$ models instead of a single one drastically increases the training time and memory cost, which may be intractable for large language models. Second, the optimization may be subject to collapse, where the same models are chosen as winners through the iterations, leaving the other models untrained. In the next section, we describe how we solve these issues with our approach LoRA-MCL.
@@ 3.3 LoRA-MCL method
Multiple choice learning typically alleviates the high training cost issue of $K$ models by training a single model with several heads lee2016stochastic ; lee2017confident . However, we argue that such an approach is not well-suited for fine-tuning language models. First, heads of most language models are quite large (for example in Qwen2-Audio chu2024qwen2 the lm_head has $d\times|\mathcal{V}|=4096\times 156032\simeq 640\mathrm{M}$ params), and standard MCL would not scale easily with the number of heads. Second, the initialization of the heads poses several challenges. Initializing each head with the parameters of the head of the pretrained model requires special care, as the collapse of the predictions is likely given the very similar hypotheses (same parameters). A complete re-initialization of the heads is detrimental to performance, as numerous training iterations would be necessary to reach the same level of knowledge as in the pretrained head. For these reasons, we consider a Low Rank Adapter (LoRA) approach hu2022lora due to its excellent trade-off between performance and computational requirements, as well as its wide adoption in the context of large model fine-tuning.
@@
$\theta_{k}=\theta\cup\left\{(A_{\ell}^{k},B_{\ell}^{k})\;|\;\ell=1,\dots,L\right\}\;,$ (4)
be the set of parameters that are involved in hypothesis $k$, with $L$ being the total number of layers where LoRA is used.
Training in the WTA fashion involves computing $p(x\,|\,c;\theta_{k})$ for $k=1,\dots,K$. To avoid collapse into a single head, we may use a (relaxed) winner-takes-all loss as a training objective:
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{c,x}\Big{[}\sum_{k=1}^{K}q_{k}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (5)
where $\{q_{k}\}$ is a set of coefficients that sum to $1$ and that give more weight to the winner index $q_{k^{\star}}$ than to the others, but still provides gradient weight to non-winner heads in order to avoid collapse. In practice, we use $q_{k^{\star}}=1-\varepsilon$ and $q_{k}=\frac{\varepsilon}{K-1}$ for $k\neq k^{\star}$ as in rupprecht2017learning .
@@ 3.4 Accelerating LoRA-MCL training with parallelization over the hypotheses
To alleviate this issue, we propose the following methodology. Let $\mathbf{x}\in\mathbb{R}^{b\times\mathcal{T}\times d}$ denote the input of the transformer architecture, where $b$ is the batch size, $\mathcal{T}\triangleq T+\tau$ is the total sequence length, and $d$ is the number of features. We duplicate $\mathbf{x}$, $K$ times along the batch size dimension to get $\mathbf{x}^{\prime}=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{bK\times\mathcal{T}\times d}$.
LoRA units at layer $\ell$ are then computed as:
$\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]\leftarrow\left[\begin{array}[]{cccc}B^{1}_{\ell}A^{1}_{\ell}&0&0&0\\ 0&B^{2}_{\ell}A^{2}_{\ell}&0&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&B^{K}_{\ell}A^{K}_{\ell}\end{array}\right]\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]+\left[\begin{array}[]{c}f_{\theta}^{\ell}(\mathbf{x}_{1})\\ f_{\theta}^{\ell}(\mathbf{x}_{2})\\ \vdots\\ f_{\theta}^{\ell}(\mathbf{x}_{K})\end{array}\right]\,,$ (6)
where $f_{\theta}^{\ell}$ is the base model (whose parameters remain frozen during training). In practice, this computation is conducted using a Group Convolution operation.111 This can be done by first reshaping $\mathbf{x}^{\prime}$ to shape $(b\times\mathcal{T}\times Kd)$ and applying a nn.Conv1d($Kd$, $Kd$, kernel_size=1, groups=$K$) following the PyTorch layer implementation, reshaping back $\mathbf{x}^{\prime}$ to shape $(Kb\times\mathcal{T}\times d)$ before adding the base model output, and then repeating at the next LoRA unit. While this duplication virtually multiplies the batch size by $K$, the memory overhead remains manageable, assuming that $r\ll d$. The next section is dedicated to the inner properties of the proposed approach.

### v2

# Multiple Choice Learning of Low-Rank Adapters for Language Modeling

@@ Abstract
We propose LoRA-MCL, a training scheme that extends next-token prediction in language models with a method designed to decode diverse, plausible sentence continuations at inference time. Traditional language modeling is an intrinsically ill-posed problem: given a context, multiple “futures” may be equally plausible. Our approach leverages Multiple Choice Learning (MCL) and the Winner-Takes-All loss to efficiently handle ambiguity through Low-Rank Adaptation. We provide a theoretical interpretation of applying MCL to language modeling, assuming the data is generated from a mixture of distributions. We illustrate the proposed approach using mixtures of Markov chains. We then demonstrate with experiments on visual and audio captioning, as well as machine translation, that our method achieves high diversity and relevance in generated outputs. The accompanying code and a general-purpose package for applying LoRA-MCL to a wide range of language models are made available. [^1] https://github.com/Victorletzelter/LoRA-MCL
Winner-takes-all, Diversity, Ambiguity, Multiple Choice Learning, Language Modeling, Low Rank Adapters
@@ 1 Introduction
Predicting what a person will say next or describing the content of an audio or visual scene with text is difficult, if not impossible, to do with perfect accuracy. When the context is not informative enough, external factors may lead to different scenarios or modes of plausible text continuations (Yang et al., 2018; Dieng et al., 2020; Mei et al., 2022b). In such ambiguous tasks, the conditional distribution over the space of output sentences given the input context may be multi-modal due to the underlying inherent uncertainty (Malinin and Gales, 2020).
@@
Initially proposed for text processing, transformer-based auto-regressive language models (Radford et al., 2018, 2019) have quickly become a general framework for modeling streams of tokens, which can also represent, for instance, images and audio signals (Touvron et al., 2023; Chu et al., 2024; Liu et al., 2023). Such models are trained as next-token predictors and allow, by nature, for addressing this uncertainty, by generating plausible output sentences through a wide body of maximum a posteriori (MAP) or sampling-based decoding approaches (Welleck et al., 2024). While sampling allows exploration and diversity, it may lead to unreliable responses and requires truncation to avoid unexpected answers, partly due to the overestimation of low-probability tokens (Zhang et al., 2021; Hewitt et al., 2022). When seeking reliable and expected answers, MAP estimation techniques, like Beam Search (Lowerre, 1976), look for sentences that maximize the model’s likelihood. However, these alternatives have drawbacks as they may lack diversity, be prone to repetition loops (Keskar et al., 2019), and may sound unnatural (Holtzman et al., 2019). Some approaches, e.g., Diverse Beam Search (Vijayakumar et al., 2018), were therefore proposed to artificially increase the diversity at inference, e.g., through a diversity penalty parameter $\lambda$, to find a tradeoff between generation quality and sample diversity (Su and Collier, 2023). In contrast with these methods, our approach aims to predict diverse sentences reflecting the ambiguity of the input context.
@@
Multiple Choice Learning (MCL) (Guzman-Rivera et al., 2012; Lee et al., 2016) has emerged as a paradigm for addressing ambiguous tasks. It generally consists of a network with a shared backbone and multiple output heads. During training, it utilizes the winner-takes-all loss for adaptively updating the head that performs the best for each example. This is a competitive training scheme that specializes each model to subsets of the conditional output distribution (Rupprecht et al., 2017). In this paper, we propose incorporating this idea for language model finetuning, leveraging multiple Low-Rank Adapters (Hu et al., 2022) instead of multiple heads, which may be impractical due to computation requirements and architectural constraints. Our method natively generates diverse and plausible sequences in a single forward pass, aiming to best approximate the conditional output distribution. Our main contributions are as follows:
@@
We propose a new paradigm that adapts MCL for token sequence modeling with LoRA-MCL, that is particularly suited for efficient finetuning of language models.
@@
We provide a theoretical analysis of our approach. Assuming the sequences are sampled from a mixture of distributions, we explain why LoRA-MCL should capture the data distribution modes, and validate the claims on Markov chains with a well-designed toy example.
We conduct extensive experiments on audio and vision captioning, as well as machine translation, showing wide applicability and an excellent diversity–quality trade-off.
@@ 2 Problem setup
$\mathcal{L}(\theta)=\mathbb{E}_{c,x}[-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)]\;,$ (1)
where $x_{<t}$ denotes the sequence of tokens prior to time $t$. Optimizing (1) is referred to as teacher-forcing (Williams and Zipser, 1989), where $p_{\theta}$ is fed with target (instead of predicted) tokens $x_{<t}$ during training (Sutskever et al., 2014). When using a transformer architecture (Vaswani et al., 2017), this is implemented via causal attention modules, which allow for computing the conditional distributions in parallel through all time steps within a single forward.
During inference, decoding methods (Welleck et al., 2024) proceed to generating sequences $\hat{x}$ from the trained model $p_{\theta}$ in an auto-regressive fashion. First, they start with a conditional distribution for the first token from the context: $p_{\theta}(x_{1}\,|\,c)$, which allows selecting $\hat{x}_{1}$. Then, for $t\geq 2$ they predict $p_{\theta}(x_{t}\,|\,\hat{x}_{<t},c)$, and select $\hat{x}_{t}$, until reaching either the sequence length limit or an end-of-sentence token. The choice of the decoding method to generate $K$ candidate sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ depends on the purpose of the task, but the general goal is $(i)$ to get highly likely sentences, i.e., ones that maximize $p_{\theta}(\hat{x}\,|\,c)$; and $(ii)$ to get diverse sentences, as can be measured by $n$-gram similarity (Ippolito et al., 2019). Although this is a widely adopted paradigm, we show next how this training and decoding pipeline can be improved: instead of artificially generating diversity at inference time, we aim at learning to predict sequences that cover well the modes of the target distribution p(x|c)p(x\,|\,c).
@@ 3.1 Motivation
In language modeling, topic models (Nigam et al., 2000; Blei et al., 2003; Dieng et al., 2020) are data-generating processes in which the ground-truth probability distribution of sequences is modeled as a mixture of latent components or topics. For example, the sequence “I am eating …” may have multiple plausible continuations, but the likelihood of each depends heavily on contextual factors such as the speaker’s location, which influences their culinary habits. Each location (or context) can thus be associated with a distinct word distribution. In topic models, data generation proceeds by first sampling a topic $z\in\mathcal{Z}$ for each sentence (usually referred to as a document in the literature), and then sampling words (or $n$-grams) from the distribution associated with that topic.
With this in mind, MLE in (1) may not be suitable (Yang et al., 2018). While MLE is effective for estimating the distribution $p(x)$, it does not capture the components when it is expressed as a mixture, i.e., $p(x)=\sum_{k}p(z_{k})p(x\,|\,z_{k})$. In such cases, MLE tends to model the aggregate rather than distinguish topic-specific distributions $p(x\,|\,z_{k})$.
3.2 Applying MCL to language modeling
Our approach is inspired by the multiple choice learning (MCL) literature (Guzman-Rivera et al., 2012; Lee et al., 2016). We propose the following training scheme, intending to enable the recovery of the different topics $z_{k}$. Instead of a single model, we consider a set of models $(\theta_{1},\dots,\theta_{K})$. Then the objective (1) is replaced by one consisting of iterating between the following two steps:
@@ 3.2 Applying MCL to language modeling
    For each training sample $(c,x)$ in the batch: Compute $p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$, and choose the best model $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$.
@@
    $\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big[\max_{k=1,\dots,K}\log p(x\,|\,c;\theta_{k})\Big]$ (2)
@@
This training procedure, similar to a hard-EM style optimization (Min et al., 2019; Wen et al., 2023), is a competitive training scheme that encourages the different models to explore different areas of the data distribution. However, it is subject to two main issues: First, using $K$ models instead of a single one drastically increases the training time and memory cost, which may be intractable for large language models (LLMs). Second, the optimization may be subject to collapse, where the same models are chosen as winners through the iterations, leaving the other models untrained. In the next section, we describe how we solve these issues with our approach LoRA-MCL.
@@ 3.3 LoRA-MCL method
Multiple choice learning typically alleviates the high training cost issue of $K$ models by training a single model with several heads (Lee et al., 2016, 2017). However, we argue that such an approach is not well-suited for fine-tuning language models. First, heads of most language models are quite large (for example, in Qwen2-Audio (Chu et al., 2024) the lm_head has $d\times|\mathcal{V}|=4096\times 156032\simeq 640\mathrm{M}$ parameters), and standard MCL would not scale easily with the number of heads. Second, the initialization of the heads poses several challenges. Initializing each head with the parameters of the head of the pretrained model requires special care, as the collapse of the predictions is likely given the very similar hypotheses (same parameters). A complete re-initialization of the heads is detrimental to performance, as numerous training iterations would be necessary to reach the same level of knowledge as in the pretrained head. These effects are empirically verified in Apx. E.6.3. For these reasons, we consider a Low Rank Adapter (LoRA) approach (Hu et al., 2022) due to its excellent trade-off between performance and computational requirements, as well as its wide adoption in the context of large model fine-tuning.
@@
$\theta_{k}=\theta\cup\left\{(A_{\ell}^{k},B_{\ell}^{k})\;|\;\ell\in\{1,\dots,L\}\right\}\;,$ (3)
be the set of parameters that are involved in hypothesis $k$, with $L$ being the total number of layers where LoRA is used. Training in the WTA fashion involves computing $p(x\,|\,c;\theta_{k})$ for $k=1,\dots,K$. To avoid situations where some heads may be under-trained, including the collapse when a single head is trained, we use the relaxation of the winner-takes-all training objective of the form
$\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big[\sum_{k=1}^{K}q_{k}\log p(x\,|\,c;\theta_{k})\Big]\;.$ (4)
where $\{q_{k}\}$ is a set of positive coefficients that sum to 1. These coefficients assign higher weight to the winning head $q_{k^{\star}}$ while still providing nonzero gradient contributions to the other heads, thereby mitigating collapse. We experimented with two relaxation techniques. First, Relaxed-WTA (Rupprecht et al., 2017) where $q_{k^{\star}}=1-\varepsilon$ and $q_{k}=\frac{\varepsilon}{K-1}$ for $k\neq k^{\star}$, with $\varepsilon>0$ a small constant. We also considered the annealed MCL method (Perera et al., 2024), which introduces a temperature parameter $\uptau$:
$q_{k}(x,c;\uptau)=\frac{p(x\,|\,c;\theta_{k})^{\frac{1}{\uptau}}}{Z_{x,c}(\uptau)}\,,\;Z_{x,c}(\uptau)=\sum_{s=1}^{K}p(x\,|\,c;\theta_{s})^{\frac{1}{\uptau}}\;.$ (5)
Here the temperature $t\mapsto\uptau(t)$ follows a decreasing schedule, typically $\uptau(t)=\uptau(0)\rho^{t}$ with $\rho<1$ and $\uptau(0)>0$. At high temperatures, training is distributed more evenly across all hypotheses, preventing collapse; as $\uptau\to 0$, the method converges to the greedy WTA regime.
[…]
Figure 1: LoRA-MCL. Components of a linear layer $\ell$ where LoRA is enabled, with context $c$ omitted. Frozen base weights $W_{\ell}$ are in blue; trainable LoRA adapters in light red. The forward pass (in gray) is computed independently for each hypothesis, where $h^{(1)},\dots,h^{(K)}$ denote the hidden states as in (34). Gradients (purple arrows) are stronger for the winning hypothesis ($k^{\star}$).
@@ 3.4 Accelerating LoRA-MCL training with parallelization over the hypotheses
A naive implementation of LoRA-MCL would require looping over the $K$ hypotheses in the batch to evaluate each candidate separately, which would drastically slow down training. To avoid this, we process all hypotheses in parallel. Specifically, given an input sequence, we duplicate it $K$ times along the batch dimension. Each copy is then passed through a LoRA-adapted transformer, but crucially, we implement this in a grouped fashion: instead of running $K$ independent forward passes, we combine them into a single batched operation where each group corresponds to one hypothesis. In practice, this can be achieved using a grouped 1D convolution (nn.Conv1d in PyTorch) with $K$ groups, so that each hypothesis uses its own LoRA weights while still sharing the frozen base model. This trick effectively multiplies the batch size by $K$ while keeping the memory overhead manageable (since the LoRA rank $r\ll d$). It removes the sequential loop, enabling efficient parallel training of all hypotheses. Details are in Apx. E.5.